OpenStudy (anonymous):

Write the expression as a single natural logarithm. 3 ln 2 + 6 ln y? can you explain how to solve it, i want to learn not just get it the easy way. thank you

jimthompson5910 (jim_thompson5910):

Use these rules y*ln(x)=ln(x^y) and ln(A)+ln(B)=ln(A*B)

jimthompson5910 (jim_thompson5910):

So y*ln(x)=ln(x^y) 3*ln(2)=ln(2^3) 3*ln(2)=ln(8)

jimthompson5910 (jim_thompson5910):

and 6*ln(y)=ln(y^6)

jimthompson5910 (jim_thompson5910):

both were found using the first rule

jimthompson5910 (jim_thompson5910):

now we go from 3*ln(2)+6*ln(y) to ln(8)+ln(y^6)

jimthompson5910 (jim_thompson5910):

now use the second rule to simplify ln(8)+ln(y^6)

OpenStudy (anonymous):

ln(8)+ln(y^6) would be ln(8y^6) ? using the equation thank you very much this helped me understand!

jimthompson5910 (jim_thompson5910):

correct, ln(8y^6) is the final answer